# The Riemann Hypothesis is false

### v44

@@
Let $\{{y\}}$ denote the fractional part of $y$. Recall that if $u$ is not an integer, then $\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi(\frac{1}{2}-\{{u\}}+O(1/X))$ as $X\rightarrow\infty$. Hence if $T\in\mathbb{Q}$ so that $T/2\pi m$ is irrational for $m\in\mathbb{N}$, we have
$\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}=\sum_{1\leq r\leq T^{1/4}}\frac{2i\sin(rT/m)}{r}=2\pi i\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O(T^{-1/4})\Big{)}.$ (30)
Plugging (30) into (29) gives
@@
$=\frac{2\pi i}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O(T^{-1/4})\Big{)}+O(T^{3/4}\log^{6}T).$ (32)
@@
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(2n/m)}$ (38)
@@
$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{H}{2\pi}+\frac{H}{2\pi}\log\pi+\frac{T}{2}+o(T)\Big{)}$ (41)
$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{T}{\pi}+\frac{T}{\pi}\log\pi+\frac{T}{2}+o(T)\Big{)}$ (42)
@@
$=\frac{2\pi i}{e}\Big{(}(1-B)\frac{T}{\pi}+\frac{T}{\pi}\log(\pi/2)+o(T)\Big{)}.$ (45)
@@
$\sum_{\frac{T}{\log^{2}T}<2n\leq T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<2n\leq T}\frac{\Lambda(2n)}{2n}+(1-B)\frac{T}{\pi}+\frac{T}{\pi}\log(\pi/2)+o(T)$ (46)
$=(1-B)\frac{T}{\pi}+\frac{T}{\pi}\log(\pi/2)+o(T)$ (47)
@@
$=(1-B+\log\pi)\frac{T}{\pi}+\frac{T}{2\pi}\log 2+o(T)$ (52)
Suppose that $\Theta<1$. Then by Lemma 1 we have
@@
$\sum_{1\leq n\leq\frac{T}{2}}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq\frac{T}{2}}(2n)^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (54)
But (54) is a contradiction to (52), thus our supposition must be false. This completes the proof. ∎
@@
From the literature on arXiv and elsewhere, it doesn’t appear like this approach on RH has been tried before. However, the idea of applying Gonek’s formula (48) in estimating sums over zeros of zeta in general, is not new. As one can confirm from the literature, it has proven to be a useful tool for such tasks. It is therefore instructive to sum (9) carefully, e.g. by not rushing to naively bound $(2n)^{it}$ by 1 for $2n$ not too close to $m$. Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sums over $2n\leq\frac{m}{1+\varepsilon}$ and $2n\geq m(1+\varepsilon)$. Note that the analogue of (48) from the Riemann explicit formula is $\sum_{T\leq\gamma_{T}\leq H}\frac{(2n)n^{\rho_{T}-1}}{\rho_{T}}\ll\frac{(\log T)(\log\log T)}{T}.$ This has no main term, so it’s not quite useful for our purposes.

### v45

@@
Let $\{{y\}}$ denote the fractional part of $y$ and $\mid\mid y\mid\mid$ denote the distance to the nearest integer to $y$. Note that if $u>0$ is not an integer, we have
$\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi\Big{(}\frac{1}{2}-\{{u\}}+O\Big{(}\frac{1}{1+||u||X}\Big{)}\Big{)}$ (30)
as $X\rightarrow\infty$. Hence if $T\in\mathbb{Q}$ so that $T/2\pi m$ is irrational for $m\in\mathbb{N}$, we have
$\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}=\sum_{1\leq r\leq T^{1/4}}\frac{2i\sin(rT/m)}{r}=2\pi i\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O\Big{(}\frac{1}{1+||T/2\pi||T^{1/4}}\Big{)}\Big{)}.$ (31)
Actually for $20<||u||=u<1/$ 2, the error term in (30) is $\ll\frac{1}{u(X+\frac{1}{2})}$ and for $||u||\geq 1/2$ it remains as it is 222see e.g. https://math.stackexchange.com/questions/3331510/error-estimation-for-the-fourier-series-of-the-fractional-part-of-x. Thus, applying (31) into (30) with $u=T/2\pi m$ and splitting the error term for $20<||T/2\pi m||<1/$ 2 and $||T/2\pi m||\geq 1/2$ gives
@@
$=\frac{2\pi i}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}-\{{T/2\pi m\}}\Big{)}+O(T^{3/4}\log^{6}T).$ (33)
@@
$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{H}{2\pi}-\frac{H}{2\pi}\log\pi+\frac{T}{2}+o(T)\Big{)}$ (42)
$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{T}{\pi}-\frac{T}{\pi}\log\pi+\frac{T}{2}+o(T)\Big{)}$ (43)
@@
$=\frac{2\pi i}{e}\Big{(}(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(\pi/2)+o(T)\Big{)}.$ (46)
@@
$\sum_{\frac{T}{\log^{2}T}<2n\leq T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<2n\leq T}\frac{\Lambda(2n)}{2n}+(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(\pi/2)+o(T)$ (47)
$=(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(\pi/2)+o(T).$ (48)
@@
$=(1-B+\log\pi)\frac{T}{2\pi}-\frac{T}{\pi}\log 2+o(T)$ (53)
for the sequence of $T,H\in[2T,2T+1]$ that satisfy (1). Suppose that $\Theta<1$. Then by Lemma 1 we have
@@
$\sum_{1\leq n\leq\frac{T}{2}}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq\frac{T}{2}}(2n)^{\rho_{T}-1}\ll T^{\Theta}\log T$ (55)
uniformly for $T\geq T_{0}$. But (55) is a contradiction to (53), thus our supposition must be false. This completes the proof. ∎
@@
From the literature on arXiv and elsewhere, it doesn’t appear like this approach on RH has been tried before. However, the idea of applying Gonek’s formula (49) in estimating sums over zeros of zeta in general, is not new. As one can confirm from the literature, it has proven to be a useful tool for such tasks. It is therefore instructive to sum (9) carefully, e.g. by not rushing to naively bound $(2n)^{it}$ by 1 for $2n$ not too close to $m$. Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sums over $2n\leq\frac{m}{1+\varepsilon}$ and $2n\geq m(1+\varepsilon)$. Note that the analogue of (49) from the Riemann explicit formula is $\sum_{T\leq\gamma_{T}\leq H}\frac{(2n)^{\rho_{T}-1}}{\rho_{T}}\ll\frac{(\log T)(\log\log T)}{T}.$ This has no main term, so it’s not quite as elegant/useful as (49).